OpenStudy (anonymous):

Solve by completing the square: -3x^2-6x-5=0

OpenStudy (anonymous):

I seem to have gone wrong somewhere in my working I checked my answer on wolframalpha but it is a little off...

jimthompson5910 (jim_thompson5910):

Can you post your work so we can have a look at it to see where you went wrong?

mathslover (mathslover):

\[-3x^2-6x-5=0\] solve this with quadratic equation : \[-b \pm \sqrt{b^2-4ac}/2a\]

mathslover (mathslover):

wait i am posting the image

OpenStudy (anonymous):

I see the mistake I made now but it's still not the same answer I got \[x=1 \pm \frac{2}{3}i\]

jimthompson5910 (jim_thompson5910):

-3x^2-6x-5=0 -3x^2-6x=5 -3(x^2+2x)=5 -3(x^2+2x+1-1)=5 -3((x^2+2x+1)-1)=5 -3((x+1)^2-1)=5 -3(x+1)^2-3(-1)=5 -3(x+1)^2+3=5 -3(x+1)^2+3-5=0 -3(x+1)^2-2=0

jimthompson5910 (jim_thompson5910):

I'll let you take it from here

OpenStudy (anonymous):

I see where I went wrong. Thanks everyone!!!